# IUT-Plug: A Plug-in tool for Interleaved Image-Text Generation

### v2

@@ Abstract
Existing vision language models (VLMs), including GPT-4 and DALL·E, often struggle to preserve logic, object identity, and style in multimodal image-text generation. This limitation significantly hinders the generalization capability of VLMs in complex image-text input-output scenarios. To address this issue, we propose IUT-Plug, a module grounded in an Image Understanding Tree (IUT), which enhances existing interleaved VLMs through explicit structured reasoning, thereby mitigating context drift in logic, entity identity, and style. The proposed framework operates in two stages. (1) A dynamic IUT-Plug extraction module parses visual scenes into hierarchical symbolic structures. (2) A coordinated narrative-flow and image synthesis mechanism ensures cross-modal consistency. To evaluate our approach, we construct a novel benchmark based on 3,000 real human-generated question-answer pairs over fine-tuned large models, introducing a dynamic evaluation protocol for quantifying context drift in interleaved VLMs. Experimental results demonstrate that IUT-Plug not only improves accuracy on established benchmarks but also effectively alleviates the three critical forms of context drift across diverse multimodal question answering (QA) scenarios.
@@ 1 Introduction
Despite remarkable advancements, modern Vision Language Models (VLMs) suffer from a fundamental limitation known as multimodal context drift. Multimodal context drift refers to a progressive loss of cross-modal consistency during extended interleaved image-text interactions. Although VLMs excel at interpreting static image-text pairs (Rombach et al., 2022; Ramesh et al., 2022), they often fail to maintain coherence across multi-turn sequences. Specifically, (1) Logic drift occurs when the joint semantic content derived from the input image and accompanying text is not faithfully reflected in either the textual answer or the subsequently generated image, leading to contradictions between perception, instruction, and output. (2) Entity identity drift manifests when concrete referents—such as named characters or objects present in the input image or text—are misidentified, lose their attributes, or disappear entirely in later generations. (3) Style drift happens when distinctive visual characteristics of the input image, including artistic medium, color palette, lighting, or composition, are not preserved in the output image despite being explicitly conditioned on that input (Malakouti and Kovashka, 2025; Said et al., 2025; Park et al., 2021). These three forms of drift reveal a systemic failure in current VLMs to maintain a unified, persistent representation of the multimodal scene across turns (Bougzime et al., 2025; Marcus, 2020).
To bridge this gap, we propose IUT-Plug, a lightweight and model-agnostic plug-in module. IUT-Plug is built around Image Understanding Trees (IUTs), which are hierarchical symbolic structures. These structures explicitly represent the entities in a visual scene, their attributes, and the relationships among them. The module integrates seamlessly with existing large vision-language models without requiring costly end-to-end retraining (Li et al., 2024; Chen et al., 2024). It operates within the neuro-symbolic paradigm (d’Avila Garcez and Lamb, 2020; Sarker et al., 2021). By design, IUT-Plug separates the perceptual strength of neural networks from the logical precision of symbolic reasoning. It uses the IUT as a factual scaffold to guide both text and image generation (Ferreira et al., 2023; Pan et al., 2024). This combination allows the system to maintain consistency while preserving the flexibility of modern generative models.
IUT-Plug bridges perception and generation. In the perception phase, the input image is parsed into an Image Understanding Tree (IUT), which represents entities, their attributes, and their relationships in a structured symbolic format. As new interleaved image-text instructions arrive, this state is incrementally updated to reflect evolving semantics. The refined IUT then actively guides both the language response and the prompt for the downstream text-to-image generator. By anchoring the entire generation process in this symbolic scaffold, IUT-Plug ensures that logical coherence, entity identity, and visual style are faithfully propagated from the input to all subsequent outputs. This design directly addresses the three core forms of multimodal context drift that undermine current interleaved VLM pipelines.
Our principal contributions are as follows. (1) We propose IUT-Plug, a lightweight and model-agnostic plug-in module. It is designed to mitigate multimodal context drift in interleaved image-text generation. IUT-Plug works by constructing an Image Understanding Tree (IUT). The IUT is a dynamic, hierarchical symbolic representation of the input image. This representation explicitly captures logic, entity identity, and visual style. IUT-Plug uses the IUT to guide both the language response and the prompt for the downstream text-to-image generator. This ensures that critical contextual information is preserved across modalities. Importantly, our approach requires no modification to the base VLM or text-to-image model. (2) We also introduce an evaluation framework. For each input–output instance, the framework dynamically generates natural-language evaluation criteria. These criteria probe consistency along the three dimensions of context drift: logic, entity identity, and style. A fine-tuned vision-language model then scores each criterion. The evaluator is trained on 3,000 human-annotated samples. It achieves 87.6% agreement with human judgment. (3) Using this framework, we demonstrate that IUT-Plug consistently improves contextual fidelity. The gains hold across diverse VLM pipelines and benchmarks. Our results establish a new standard for measuring and achieving compositional consistency in interleaved multimodal generation.

### v3

# Bridging the Consistency Gap: Explicit Structured Memory for Interleaved Image-Text Generation

@@ Abstract
Existing Vision Language Models (VLMs) often struggle to preserve logic, entity identity, and artistic style during extended, interleaved image-text interactions. We identify this limitation as ”Multimodal Context Drift”, which stems from the inherent tendency of implicit neural representations to decay or become entangled over long sequences. To bridge this gap, we propose IUT-Plug, a model-agnostic Neuro-Symbolic Structured State Tracking mechanism. Unlike purely neural approaches that rely on transient attention maps, IUT-Plug introduces the Image Understanding Tree (IUT) as an explicit, persistent memory module. The framework operates by (1) parsing visual scenes into hierarchical symbolic structures (entities, attributes, and relationships); (2) performing incremental state updates to logically lock invariant properties while modifying changing elements; and (3) guiding generation through topological constraints. We evaluate our approach on a novel benchmark comprising 3,000 human-annotated samples. Experimental results demonstrate that IUT-Plug effectively mitigates context drift, achieving significantly higher consistency scores compared to unstructured text-prompting baselines. This confirms that explicit symbolic grounding is essential for maintaining robust long-horizon consistency in multimodal generation.
@@ 1 Introduction
Despite the significant progress achieved by Multimodal Large Language Models (MLLMs) in single-turn image-text understanding and generation tasks (Rombach et al., 2022; Ramesh et al., 2022), existing mainstream architectures encounter severe challenges when handling long-horizon, multi-turn interleaved image-text interactions. We identify this critical issue as Multimodal Context Drift. Pure text Large Language Models (LLMs) maintain semantic consistency effectively via the Key-Value (KV) Cache mechanism, which stores vector representations of preceding tokens to avoid redundant computations during the generation of subsequent tokens. In contrast, visual understanding and generation models frequently suffer from catastrophic forgetting in multi-turn dialogues. For most text-to-image or text-to-video models, subsequent generation turns are essentially independent processes. Mainstream visual architectures lack a persistent memory mechanism analogous to the textual KV Cache to preserve the pixel-level states or high-level semantic features of previous images. Consequently, as the number of interactions increases, the model progressively loses its ability to enforce constraints regarding initial visual entities, artistic styles, and logical relationships.
Current visual understanding and generation models trained under the End-to-End paradigm exhibit inherent limitations. They primarily rely on implicit neural representations (d’Avila Garcez and Lamb, 2020; Sarker et al., 2021) and transient attention mechanisms to propagate visual information, a reliance that directly contributes to context drift. In complex, long-horizon interactions, high-dimensional visual features are prone to entanglement with textual features or dilution by noise. For instance, when a user demands that the cat ”jump,” the model may correctly generate the action but fail to retain the cat’s fur pattern” or “background style” defined in the initial turn. Mere scaling of model parameters is insufficient to thoroughly resolve this issue of fine-grained state tracking (Bougzime et al., 2025; Marcus, 2020).
To bridge the aforementioned gap, we introduce the Image Understanding Tree (IUT) as the core data structure. Unlike previous approaches that attempt to improve consistency through full-model fine-tuning, IUT-Plug is centered on the principles of explicit disentanglement and state locking. The IUT is not a mere image caption; rather, it is a hierarchical symbolic state representation that dynamically stores and updates entity attributes, spatial relationships, and global styles within the scene. As illustrated in Figure  2: (1) Structured Perception: This leverages a Vision-Language Model (VLM) to parse unstructured visual input into a structured IUT state tree. (2) Incremental State Tracking: During multi-turn interactions, this module functions as an “external memory.” It updates only the nodes in the IUT that have changed according to user instructions (e.g., actions that alter an entity) while strictly locking unchanged attributes (such as entity IDs and styles), thereby mechanically blocking information drift. (3) Constrained Generation: The updated IUT is serialized into a prompt with topological constraints, which guides the T2I (Text-to-Image) model in generation.
Our approach not only obviates the need for computationally expensive full-parameter fine-tuning but also provides an interpretable and intervenable intermediate layer (Li et al., 2024; Chen et al., 2024). To validate the efficacy of IUT-Plug, we constructed a dynamic evaluation benchmark comprising 3,000 real-world user interaction samples. Experimental results demonstrate that IUT-Plug achieves significant improvements in Entity Consistency and Logical Consistency compared to strong baselines that rely solely on textual context memory (Structured Text Prompting). These findings confirm that explicit structured memory is pivotal for mitigating semantic drift in long-horizon multimodal generation.